# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 14.4 Demazure functors
Follows from rationality of Demazure resolutions applied to the longest word $w_{0}\in W$; more generally, ${d_{w}}_{\ast}\mathcal{O}_{Z_{w}}=\mathcal{O}_{X_{w}}$, where $d_{w}:Z_{w}\rightarrow X_{w}$ is a Demazure resolution of the Schubert variety $X_{w}$ and $w\in W$ is a reduced word, [11]. ∎
@@
Let $\omega_{i}$ be a fundamental weight. The space of global sections $\Gamma(X,\mathcal{L}({\omega_{i}}))$ is isomorphic to $\nabla_{\omega_{i}}$ (the induced module). Let $\omega_{X}$ (resp., $\omega_{Y_{i}}$) denote the canonical line bundle on $X$ (resp., on $Y_{i}$). Then ${\pi_{\alpha_{i}}}_{\ast}\omega_{X}=\omega_{Y_{i}}[-1]$. The relative canonical sheaf $\omega_{X/Y_{i}}$ is isomorphic to $\mathcal{L}(-\alpha_{i})$ [16, Lemme 6], and hence ${\pi_{i}}_{\ast}\mathcal{L}({-\alpha_{i}})=\mathcal{O}_{Y_{i}}[-1]$. More generally, given an object $\mathcal{F}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$, Grothendieck duality for the morphism ${\pi_{\alpha_{i}}}$ gives an isomorphism ${\pi_{\alpha_{i}}}_{\ast}(\mathcal{F}\otimes\mathcal{L}(-\alpha_{i}))[1]={\pi_{\alpha_{i}}}_{\ast}(\mathcal{F}^{\vee})^{\vee}$ where the inner $(-)^{\vee}$ is the duality $R{\mathcal{H}om}_{X}(\mathcal{F},\mathcal{O}_{X})$ on $X$ and the outer $(-)^{\vee}$ is the duality $R{\mathcal{H}om}_{Y_{i}}(\mathcal{F},\mathcal{O}_{Y_{i}})$ on $Y_{i}$. Since the pullback ${\pi_{\alpha_{i}}^{\ast}}$ commutes with $(-)^{\vee}$, we obtain an isomorphism ${\sf D}_{i}(\mathcal{F}\otimes\mathcal{L}(-\alpha_{i}))[1]={\sf D}_{i}(\mathcal{F}^{\vee})^{\vee}$.
Given a line bundle $\mathcal{L}(\chi),\chi\in X({T})$ and a minimal parabolic ${P}_{\alpha}$ generated by $B$ and $U_{\alpha}$, the degree of the restriction of $\mathcal{L}(\chi)$ to fibres of the projection $\pi_{\alpha}:{G}/{B}\rightarrow{G}/{P}_{\alpha}$ is equal to $\langle\chi,\alpha^{\vee}\rangle$, [15, Lemme 2].
Given a fundamental weight $\omega_{i}$, there is a short exact sequence [16, Section 3]:
@@
14.5 Demazure resolutions
Let $w\in W$ and $s_{\alpha_{i}}s_{\alpha_{i_{2}}}\dots s_{\alpha_{i_{k}}}$ be a reduced expression of $w$. Associated to $s_{\alpha_{i}}s_{\alpha_{i_{2}}}\dots s_{\alpha_{i_{k}}}$ is the Bott-Samelson-Demazure-Hansen resolution $Z_{w}$ of the Schubert variety $X_{w}$ associated to $w$, [7, Section 2.2] and [47, Section 1.2]. Both $X_{w}$ and $Z_{w}$ are acted upon by $B$ and one can form $G$-varieties by setting ${\mathcal{X}}_{w}=G\times^{B}X_{w}\subset X\times X$ and ${\mathcal{Z}}_{w}=G\times^{B}Z_{w}$. Then ${\mathcal{Z}}_{w}$ is a resolution of singularities of ${\mathcal{X}}_{w}$. Note that the variety ${\mathcal{Z}}_{w}$ depends on the choice of a reduced expression of $w$.
Theorem 14.9.
[47, Corollary 2.2.7]
${\sf D}_{w}=\Phi_{\mathcal{O}_{{\mathcal{X}}_{w}}}={q_{w}}_{\ast}p_{w}^{\ast}$ (48)
Proof.
14.6 Braid relations for Demazure functors
Let $(W,S=\{s_{\alpha_{1}},\dots,s_{\alpha_{r}}\})$ be the Coxeter presentation of the Weyl group of a simple simply connected group $G$ of rank $r$. The basic relations are $(s_{\alpha_{i}}s_{\alpha_{j}})^{m_{ij}}=1$ where $m_{ij}\in\{2,3,4,6\}$.
The braid relations hold for Demazure operators ${\sf D}_{i}$’s:
$\underbrace{{\sf D}_{i}{\sf D}_{j}{\sf D}_{i}\dots}_{m_{ij}}=\underbrace{{\sf D}_{j}{\sf D}_{i}{\sf D}_{j}\dots}_{m_{ij}}$ (49)
(i.e., each side is a product of $m_{ij}$ Demazure operators). Their validity follows from Theorem 14.9.
In type ${\sf A}_{n}$ all the exponents $m_{ij}\leq 3$ and the braid relations simplify to
${\sf D}_{i}{\sf D}_{j}{\sf D}_{i}={\sf D}_{j}{\sf D}_{i}{\sf D}_{j}$ (50)
if $\langle\alpha_{i},\alpha_{j}^{\vee}\rangle=-1$; otherwise ${\sf D}_{i}$ and ${\sf D}_{j}$ commute:
${\sf D}_{i}{\sf D}_{j}={\sf D}_{j}{\sf D}_{i}$ (51)
if $\langle\alpha_{i},\alpha_{j}^{\vee}\rangle=0$.
@@ 14.8 An exact triangle
Let $\mathcal{E}$ be a vector bundle of rank $r$ on $X$, and consider the associated projective bundle $\pi:\mathbb{P}(\mathcal{E})\rightarrow X$. Denote $\mathcal{O}_{\pi}(-1)$ the line bundle on $\mathbb{P}(\mathcal{E})$ of relative degree $-1$, such that $\pi_{\ast}\mathcal{O}_{\pi}(1)=\mathcal{E}^{\ast}$. One has [26, Corollary 8.36]:
@@
$\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathbb{P}(\mathcal{E}))=\langle\pi^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)\otimes\mathcal{O}_{\pi}(-r+1),\dots,\pi^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)\otimes\mathcal{O}_{\pi}(-1),\pi^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)\rangle.$ (53)
@@
Recall the formula for the functor $Q(-)$ (Section 2.3 and [47, Definition 7.2.1]):
$Q:={{\mathrm{res}}}_{B}^{P_{\alpha_{1}}}({\mathbb{Z}}_{-\rho}\otimes{{\mathrm{ind}}}_{B}^{P_{\alpha_{1}}}({\mathbb{Z}}_{s_{\alpha_{1}}(\rho)}\otimes_{\mathbb{Z}}(-))\dots{\mathrm{res}}_{B}^{P_{\alpha_{n}}}({\mathbb{Z}}_{-\rho}\otimes{{\mathrm{ind}}}_{B}^{P_{\alpha_{n}}}({\mathbb{Z}}_{s_{\alpha_{n}}(\rho)}\otimes_{\mathbb{Z}}(-))$ (55)
The cone ${\sf D}_{i}(-\otimes\mathcal{L}(-\omega_{i}))\otimes\mathcal{L}(\omega_{i}-\alpha_{i}))$ in (54) of Proposition 14.12 is the derived version of the factor ${{\mathrm{res}}}_{\bf B}^{{\bf P}_{\alpha_{i}}}({\mathbb{Z}}_{-\rho}\otimes{{\mathrm{ind}}}_{\bf B}^{{\bf P}_{\alpha_{1}}}({\mathbb{Z}}_{s_{\alpha_{i}}(\rho)}\otimes_{\mathbb{Z}}(-))$ of (55): note that $\langle s_{\alpha_{i}}(\rho),\alpha_{i}^{\vee}\rangle=\langle-\omega_{i},\alpha_{i}^{\vee}\rangle=-1$ and $\langle-\rho,\alpha_{i}^{\vee}\rangle=\langle\omega_{i}-\alpha_{i},\alpha_{i}^{\vee}\rangle=-1$. Iterating the cones, we get a right derived functor of $Q$.
@@ 14.9 More on Demazure functors
Let $G$ be a simple simply connected algebraic group and $\omega_{1},\dots,\omega_{r}$ be the fundamental weights of ${G}$. Given a fundamental weight $\omega_{i}$, define the vector bundle $\Psi_{1}^{\omega_{i}}$ as the kernel of the canonical surjective morphism $\nabla_{\omega_{i}}\otimes\mathcal{L}(0)\rightarrow\mathcal{L}(\omega_{i})$. Thus, there is the short exact sequence:
@@
Let the Dynkin type of $G$ be ${\sf A}_{n}$. Then
@@
The above Proposition 14.15 can easily be generalized to all simple groups, but the formulae for ${\sf D}_{k}((\Psi_{1}^{\omega_{k}})^{\ast}\otimes\mathcal{L}(-\omega_{k}))$ will then depend on the fundamental weight $\omega_{k}$. We record the examples of ${B}_{2}$ and of ${G}_{2}$ that we will need later. Let $\alpha,\beta$ be the simple roots with $\alpha$ being the short root. Then if $G$ is of type ${B}_{2}$, we have:
@@
Consider ${\sf D}_{w}\mathcal{L}(\omega_{i})$ which a priori belongs to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{\geq 0}(X)$, the functor ${\sf D}_{w}$ being a right derived functor. In fact, ${\sf D}_{w}\mathcal{L}(\omega_{i})$ is a coherent sheaf, i.e. a pure object of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$: this follows from Definition 1, Proposition 14.5 and the higher cohomology vanishing ${\rm H}^{k}(X_{w},\mathcal{L}(\omega_{i}))=0$ for $k>0$. Moreover, ${\sf D}_{w}\mathcal{L}(\omega_{i})$ is locally free, being a ${G}$-equivariant coherent sheaf on $X$.
The space of global sections of $\mathcal{L}(\omega_{i})$ is also isomorphic to $\nabla_{\omega_{i}}$ and remains to show that the morphism $\nabla_{\omega_{i}}\otimes\mathcal{L}(0)\rightarrow{\sf D}_{w}\mathcal{L}(\omega_{i})$ is surjective; this also follows from the surjectivity of the above restriction map on the global sections $\nabla_{\omega_{i}}\rightarrow{\rm H}^{0}(X_{w},\mathcal{L}(\omega_{i}))$ (cf. [47, Proposition A.2.6]) and applying Proposition 14.5 to that surjection. Now ${\sf D}_{w}\Psi_{1}^{\omega_{i}}$ is the kernel of a surjective map of vector bundles, hence is also a vector bundle. ∎
@@
Let $\omega_{i},\omega_{j}$ be two distinct fundamental weights of $G$. Then
@@
Let the rank of group $G$ be equal to 2, and let ${P}_{\alpha_{i}}$ denote the minimal parabolic generated by ${B}$ and $U_{\alpha_{i}}$ for a simple root $\alpha_{i},i=1,2$. Consider the subcategory $\pi_{\alpha_{1}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{1})\diamond\pi_{\alpha_{2}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{2})$, i.e. the minimal thick strict full triangulated subcategory containing both subcategories $\pi_{\alpha_{i}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{i})$ for $i=1,2$. Then
@@
Indeed, having established the above equality, we obtain (72) by tensoring (73) with $\mathcal{L}(\rho)$ and mutating the subcategory $\langle\mathcal{L}(\rho)\rangle$ to the left past its right orthogonal. By [32, Lemma 2.11], the mutation is given by the Serre functor which is isomorphic to tensoring with $\omega_{X}=\mathcal{L}(-2\rho)$.
@@
Since $\mathcal{F}$ is arbitrary and the functor ${\pi_{\alpha_{i}}}_{\ast}$ is surjective, it follows that ${\pi_{\alpha_{i}}}_{\ast}(\mathcal{G}\otimes\mathcal{L}({-\rho}))=0$. This implies $\mathcal{G}\in\pi_{\alpha_{i}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{i})$, the line bundle $\mathcal{L}({-\rho})$ having degree $-1$ along $\pi_{\alpha}$. It follows that $\sf D_{\alpha_{i}}\mathcal{G}=\mathcal{G}$ for $i=1,2$ for such an object $\mathcal{G}$. Choosing a reduced expression for the longest element $w_{0}$, we obtain $\sf D_{w_{0}}\mathcal{G}=\mathcal{G}$, and hence $\pi^{\ast}\pi_{\ast}\mathcal{G}=\mathcal{G}$, since $\sf D_{w_{0}}=\pi^{\ast}\pi_{\ast}$ by Theorem 14.6. Now $\pi^{\ast}\pi_{\ast}\mathcal{G}=\mathcal{L}(0)\otimes\pi_{\ast}\mathcal{G}=\mathcal{L}(0)\otimes\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{G})\in\langle\mathcal{L}(0)\rangle$, and the statement follows. ∎
@@
Let the rank of group $G$ be equal to 2, and $X$ the flag variety of $G$. Then $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$ has a semiorthogonal decomposition:

### v3

@@ 14.5 Demazure resolutions
Setting $w=w_{0}$ in Theorem 14.6 gives the result. ∎
@@
Let $\omega_{i}$ be a fundamental weight. The space of global sections $\Gamma(X,\mathcal{L}({\omega_{i}}))$ is isomorphic to $\nabla_{\omega_{i}}$ (the induced module). Let $\omega_{X}$ (resp., $\omega_{Y_{i}}$) denote the canonical line bundle on $X$ (resp., on $Y_{i}$). Then ${\pi_{\alpha_{i}}}_{\ast}\omega_{X}=\omega_{Y_{i}}[-1]$. The relative canonical sheaf $\omega_{X/Y_{i}}$ is isomorphic to $\mathcal{L}(-\alpha_{i})$ [19, Lemme 6], and hence ${\pi_{i}}_{\ast}\mathcal{L}({-\alpha_{i}})=\mathcal{O}_{Y_{i}}[-1]$. More generally, given an object $\mathcal{F}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$, Grothendieck duality for the morphism ${\pi_{\alpha_{i}}}$ gives an isomorphism ${\pi_{\alpha_{i}}}_{\ast}(\mathcal{F}\otimes\mathcal{L}(-\alpha_{i}))[1]={\pi_{\alpha_{i}}}_{\ast}(\mathcal{F}^{\vee})^{\vee}$ where the inner $(-)^{\vee}$ is the duality $R{\mathcal{H}om}_{X}(\mathcal{F},\mathcal{O}_{X})$ on $X$ and the outer $(-)^{\vee}$ is the duality $R{\mathcal{H}om}_{Y_{i}}(\mathcal{F},\mathcal{O}_{Y_{i}})$ on $Y_{i}$. Since the pullback ${\pi_{\alpha_{i}}^{\ast}}$ commutes with $(-)^{\vee}$, we obtain an isomorphism ${\sf D}_{i}(\mathcal{F}\otimes\mathcal{L}(-\alpha_{i}))[1]={\sf D}_{i}(\mathcal{F}^{\vee})^{\vee}$.
Given a line bundle $\mathcal{L}(\chi),\chi\in X({\mathbf{T}})$ and a minimal parabolic ${\mathbf{P}}_{\alpha}$ generated by $\mathbf{B}$ and $\mathbf{U}_{\alpha}$, the degree of the restriction of $\mathcal{L}(\chi)$ to fibres of the projection $\pi_{\alpha}:{\mathbf{G}}/{\mathbf{B}}\rightarrow{\mathbf{G}}/{\mathbf{P}}_{\alpha}$ is equal to $\langle\chi,\alpha^{\vee}\rangle$, [18, Lemme 2].
Given a fundamental weight $\omega_{i}$, there is a short exact sequence [19, Section 3]:
@@ 14.7 An exact triangle
Let $S$ be a smooth projective scheme and $\mathcal{E}$ be a vector bundle of rank $r$ on $S$. Consider the associated projective bundle $\pi:\mathbb{P}(\mathcal{E})\rightarrow S$. Denote $\mathcal{O}_{\pi}(-1)$ the line bundle on $\mathbb{P}(\mathcal{E})$ of relative degree $-1$, such that $\pi_{\ast}\mathcal{O}_{\pi}(1)=\mathcal{E}^{\ast}$. One has [29, Corollary 8.36]:
@@
$\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathbb{P}(\mathcal{E}))=\langle\pi^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(S)\otimes\mathcal{O}_{\pi}(-r+1),\dots,\pi^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(S)\otimes\mathcal{O}_{\pi}(-1),\pi^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(S)\rangle.$ (59)
@@
Note that the object $\mathcal{E}$ in the above proposition is not required to be $\mathbf{G}$-equivariant. However, if $\mathcal{F}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$ is a $\mathbf{G}$-equivariant object of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$, i.e. $\mathcal{F}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}_{\mathbf{G}}(X)=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ (and therefore $\mathcal{F}=\mathcal{L}(\sf F)$ for an object ${\sf F}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$, then the maps in the triangle (60) are $\mathbf{G}$-equivariant. More specifically, in that case triangle (60) comes from an exact triangle in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ upon applying the functor $\mathcal{L}$. Precisely, (60) is obtained from the following triangle below by applying Proposition 14.5 to it:
Proposition 14.13.
Let ${\sf F}$ be an object $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ and $\alpha_{i}\in S$ a simple root. Let ${{\sf H}}_{i}$ denote the Joseph-Demazure functor ${{\sf H}}_{s_{\alpha_{i}}}$. Then there is an exact triangle in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$
$\dots\longrightarrow{\rm R}{{\sf H}}_{i}({\sf F}\otimes{\sf k}_{-\omega_{i}})\otimes{\sf k}_{\omega_{i}-\alpha_{i}}\longrightarrow{\rm R}{{\sf H}}_{i}{\sf F}\rightarrow{\sf F}\stackrel{{\scriptstyle[1]}}{{\longrightarrow}}\dots$ (61)
@@
Recall the formula for the functor $Q(-)$ (Section 2.3 and [49, Definition 7.2.1]). Given a weight $\lambda\in X({\bf T})$, let $\lambda^{+}$ be the dominant weight in the Weyl group orbit of $\lambda$ and let $w\in W$ be minimal so that $\lambda=w\lambda^{+}$ (cf. Definition 2.2). Choose a reduced expression for $w=s_{\alpha_{1}}s_{\alpha_{2}}\dots s_{\alpha_{k}}$. For a ${\bf B}$-module $M$ and a simple root $\alpha\in S$ define the functor $F_{\alpha}(M):={{\mathrm{res}}}_{\bf B}^{{\bf P}_{\alpha}}({\mathbb{Z}}_{-\rho}\otimes{{\mathrm{ind}}}_{\bf B}^{{\bf P}_{\alpha}}({\mathbb{Z}}_{s_{\alpha}(\rho)}\otimes_{\mathbb{Z}}M))$ [^5] Recall that $\mathbf{B}$ is generated by negative roots, so the signs of $\rho$ and of $s_{\alpha_{i}}(\rho)$ in each factor are opposite to those from [49, Definition 7.2.1].. Set
$Q(\lambda)_{\mathbb{Z}}:=F_{\alpha_{1}}\dots F_{\alpha_{k}}{\mathbb{Z}}_{\lambda^{+}}.$ (62)
Then [49, Proposition 7.2.2] implies that $Q(\lambda)$ defined above is the integral form of relative Schubert modules of Definition 2.3.
We can reconcile the exact triangle (61) from Proposition 14.13 with the above description of the functor $Q$. For a simple root $\alpha_{i}$, the functor $F_{\alpha_{i}}(-)={{\mathrm{res}}}_{\bf B}^{{\bf P}_{\alpha_{i}}}({\mathbb{Z}}_{-\rho}\otimes{{\mathrm{ind}}}_{\bf B}^{{\bf P}_{\alpha_{1}}}({\mathbb{Z}}_{s_{\alpha_{i}}(\rho)}\otimes_{\mathbb{Z}}(-))$ is left exact; its right derived functor is given by the cone ${\rm R}{{\sf H}}_{i}(-\otimes{\sf k}_{-\omega_{i}})\otimes{\sf k}_{\omega_{i}-\alpha_{i}}$ in (61) of Proposition 14.13: to see it, note that $\langle s_{\alpha_{i}}(\rho),\alpha_{i}^{\vee}\rangle=\langle-\omega_{i},\alpha_{i}^{\vee}\rangle=-1$ and $\langle-\rho,\alpha_{i}^{\vee}\rangle=\langle\omega_{i}-\alpha_{i},\alpha_{i}^{\vee}\rangle=-1$. So, the triangle (61) may as well be written as follows:
$\dots\longrightarrow{\rm R}{{\sf H}}_{i}({\sf F}\otimes{\sf k}_{s_{\alpha_{i}}(\rho)})\otimes{\sf k}_{-\rho}\longrightarrow{\rm R}{{\sf H}}_{i}{\sf F}\rightarrow{\sf F}\stackrel{{\scriptstyle[1]}}{{\longrightarrow}}\dots$ (63)
Given a reduced expression $s_{\alpha_{1}}\dots s_{\alpha_{k}}$ of an element $w\in W$, there is an isomorphism of functors ${\rm R}{{\sf H}}_{w}={\rm R}{{\sf H}}_{1}\dots\rm R{{\sf H}}_{k}$. Iterating the cones in exact triangles (61) for $i=1,\dots,k$, we obtain a right derived functor of $Q$.
@@ 14.8 More on Demazure functors
Let $\mathbf{G}$ be a simple simply connected algebraic group and $\omega_{1},\dots,\omega_{r}$ be the fundamental weights of ${\mathbf{G}}$. Given a fundamental weight $\omega_{i}$, define the vector bundle $\Psi_{1}^{\omega_{i}}$ as the kernel of the canonical surjective morphism $\nabla_{\omega_{i}}\otimes\mathcal{L}(0)\rightarrow\mathcal{L}(\omega_{i})$. Thus, there is the short exact sequence:
@@
Let the Dynkin type of $\mathbf{G}$ be ${\sf A}_{n}$. Then
@@
The above Proposition 14.16 can easily be generalized to all simple groups, but the formulae for ${\sf D}_{k}((\Psi_{1}^{\omega_{k}})^{\ast}\otimes\mathcal{L}(-\omega_{k}))$ will then depend on the fundamental weight $\omega_{k}$. We record the examples of ${\mathbf{B}}_{2}$ and of ${\bf G}_{2}$ that we will need later. Let $\alpha,\beta$ be the simple roots with $\alpha$ being the short root. Then if $\mathbf{G}$ is of type ${\mathbf{B}}_{2}$, we have:
@@
Consider ${\sf D}_{w}\mathcal{L}(\omega_{i})$ which a priori belongs to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{\geq 0}(X)$, the functor ${\sf D}_{w}$ being a right derived functor. In fact, ${\sf D}_{w}\mathcal{L}(\omega_{i})$ is a coherent sheaf, i.e. a pure object of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$: this follows from Definition 2.1, Proposition 14.5 and the higher cohomology vanishing ${\rm H}^{k}(X_{w},\mathcal{L}(\omega_{i}))=0$ for $k>0$. Moreover, ${\sf D}_{w}\mathcal{L}(\omega_{i})$ is locally free, being a ${\mathbf{G}}$-equivariant coherent sheaf on $X$.
The space of global sections of $\mathcal{L}(\omega_{i})$ is also isomorphic to $\nabla_{\omega_{i}}$ and remains to show that the morphism $\nabla_{\omega_{i}}\otimes\mathcal{L}(0)\rightarrow{\sf D}_{w}\mathcal{L}(\omega_{i})$ is surjective; this also follows from the surjectivity of the above restriction map on the global sections $\nabla_{\omega_{i}}\rightarrow{\rm H}^{0}(X_{w},\mathcal{L}(\omega_{i}))$ (cf. [49, Proposition A.2.6]) and applying Proposition 14.5 to that surjection. Now ${\sf D}_{w}\Psi_{1}^{\omega_{i}}$ is the kernel of a surjective map of vector bundles, hence is also a vector bundle. ∎
@@
Let $\omega_{i},\omega_{j}$ be two distinct fundamental weights of $\mathbf{G}$. Then
@@
Let the rank of group $\mathbf{G}$ be equal to 2, and let ${\mathbf{P}}_{\alpha_{i}}$ denote the minimal parabolic generated by ${\mathbf{B}}$ and $\mathbf{U}_{\alpha_{i}}$ for a simple root $\alpha_{i},i=1,2$. Consider the subcategory $\pi_{\alpha_{1}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{1})\diamond\pi_{\alpha_{2}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{2})$, i.e. the minimal thick strict full triangulated subcategory containing both subcategories $\pi_{\alpha_{i}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{i})$ for $i=1,2$. Then
@@
Indeed, having established the above equality, we obtain (80) by tensoring (81) with $\mathcal{L}(\rho)$ and mutating the subcategory $\langle\mathcal{L}(\rho)\rangle$ to the left past its right orthogonal. By [35, Lemma 2.11], the mutation is given by the Serre functor which is isomorphic to tensoring with $\omega_{X}=\mathcal{L}(-2\rho)$.
@@
Since $\mathcal{F}$ is arbitrary and the functor ${\pi_{\alpha_{i}}}_{\ast}$ is surjective, it follows that ${\pi_{\alpha_{i}}}_{\ast}(\mathcal{G}\otimes\mathcal{L}({-\rho}))=0$. This implies $\mathcal{G}\in\pi_{\alpha_{i}}^{\ast}\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(Y_{i})$, the line bundle $\mathcal{L}({-\rho})$ having degree $-1$ along $\pi_{\alpha}$. It follows that $\sf D_{\alpha_{i}}\mathcal{G}=\mathcal{{\mathcal{G}}}$ for $i=1,2$ for such an object $\mathcal{{\mathcal{G}}}$. Choosing a reduced expression for the longest element $w_{0}$, we obtain $\sf D_{w_{0}}\mathcal{G}=\mathcal{G}$, and hence $\pi^{\ast}\pi_{\ast}\mathcal{G}=\mathcal{G}$, since $\sf D_{w_{0}}=\pi^{\ast}\pi_{\ast}$ by Theorem 14.7. Now $\pi^{\ast}\pi_{\ast}\mathcal{G}=\mathcal{L}(0)\otimes\pi_{\ast}\mathcal{G}=\mathcal{L}(0)\otimes\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{G})\in\langle\mathcal{L}(0)\rangle$, and the statement follows. ∎
@@
Let the rank of group $\mathbf{G}$ be equal to 2, and $X$ the flag variety of $\mathbf{G}$. Then $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$ has a semiorthogonal decomposition: